OpenStudy (amtran_bus):

If F(x)=2x-1/3, and F^-1 is the inverse of F, what is the value of F^-1 (3) ?

OpenStudy (whpalmer4):

Do you know how to find the inverse of F?

OpenStudy (amtran_bus):

That would be a good place to start.

OpenStudy (amtran_bus):

I need help.

OpenStudy (whpalmer4):

Okay. well, let's do a different problem as an example. \(F(x) = 3x-2\) To find the inverse function, \(F^{-1}(x)\), we solve for \(x\), then swap the variables. \[y = F(x) = 3x-2\]\[y=3x-2\]\[y+2=3x\]\[x=\frac{y+2}{3}\]Now swap the variables:\[F^{-1}(x) = \frac{x+2}{3}\]

OpenStudy (amtran_bus):

Got that.

OpenStudy (amtran_bus):

on my problem would the inverse by y=6x-3?

OpenStudy (whpalmer4):

One interesting thing about inverse functions is that if you graph a function and its inverse on the same graph, the inverse is just the function reflected across the line y = x. Here's the graph of the example I just did:

OpenStudy (whpalmer4):

So I'm not positive about your original function: is it \[F(x) = \frac{2x-1}{3}\]

OpenStudy (amtran_bus):

Sorry, yes it is.

OpenStudy (whpalmer4):

By my figuring, the inverse would then be \(y = \frac{1}{2}(1+3x)\)

OpenStudy (amtran_bus):

Do you mind showing me how?

OpenStudy (whpalmer4):

\[y = \frac{2x-1}{3}\]Solve for \(x\): \[3y = 2x-1\]\[3y+1=2x\]\[\frac{3y+1}{2} = x\]Switch variables:\[F^{-1}(x) = \frac{3x+1}{2} = \frac{1}{2}(3x+1)\]

OpenStudy (amtran_bus):

Got ya

OpenStudy (whpalmer4):

Graphed, the functions look like this:

OpenStudy (whpalmer4):

There's that reflection again...if you don't get that result, there's a mistake somewhere. So now having found the inverse function, what remains to be done is to evaluate it at \(x = 3\). Do that and tell me the answer you get...

OpenStudy (amtran_bus):

When x=3, I got 5.

OpenStudy (whpalmer4):

The olive colored line is the \(F^{-1}(x)\) graph in my second attachment. Does the graph match up with your result?

OpenStudy (amtran_bus):

Yes.

OpenStudy (amtran_bus):

One more thing. How did you get 1/2 (3x+1) from 3x+1/2?

OpenStudy (amtran_bus):

Who, nevermind. See it!

OpenStudy (whpalmer4):

Excellent! Okay, I'm going to remind you that (3x+1)/2 does not equal 3x+1/2....

OpenStudy (whpalmer4):

it's a pain putting in the extra parentheses, I agree, but probably 1/3 of the problems I help with on OS start out with me (or someone else) asking for clarification of the equation(s) for just this reason :-)

OpenStudy (amtran_bus):

Thanks. God bless.

OpenStudy (whpalmer4):

got any more like to do?

OpenStudy (amtran_bus):

No, I am reviewing for a accuplacer test to clep college algebra. I know this stuff, its just been awhile!

OpenStudy (whpalmer4):

Great! Good luck with the test...

OpenStudy (amtran_bus):

Thank you.